# Interpreting Graph Inference with Skyline Explanations

### v3

@@ VII-B Additional Experimental Results
Case Study: Edge Growing v.s. Onion-Peeling. In this case study, we compare the application scenarios of $\mathsf{ASX}$-$\mathsf{OP}$ and $\mathsf{ASX}$-$\mathsf{I}$, as shown in Figure 9. We observe that $\mathsf{ASX}$-$\mathsf{OP}$, with onion peeling, cater users’ needs who “think slow” and seeks in-depth explanations that are counterfactual. Such larger and counterfactual explanations are more likely to affect $\mathsf{GNN}$ behaviors if removed [11]. $\mathsf{ASX}$-$\mathsf{I}$, on the other hand, due to its containment of edges closer to targeted ones via edge insertions/growths, tends to produce – at its early stage – smaller, factual explanations as convincing evidences. This caters users who “think fast” and is good at quick generation of factual explanations [11]. For example, in answering “why” node $v^{\prime}_{t}$ is recognized as a “Politician” site, $G_{\zeta}^{8}$ and $G_{\zeta}^{9}$ are among skyline explanations from $\mathsf{ASX}$-$\mathsf{I}$ at early stage, revealing the fact that $v^{\prime}_{t}$ frequently interact with TV shows, Governments, and Companies, covering its ego-networks as quick facts. $G_{\zeta}^{10}$ and $G_{\zeta}^{11}$, which are reported by $\mathsf{ASX}$-$\mathsf{OP}$ in its early stage via “onion-peeling”, are relatively larger counterfactual explanations that suggest interaction patterns with more complex structures. For example, $G_{\zeta}^{11}$ includes two-hop neighbors and a triangle, which indicates the importance of the corresponding company and TV show. Most importantly, another politician’s two-hop neighbor also shares these nodes, revealing the close relationship between the two politicians.

### v4

@@ VII-B Additional Experimental Results
Case Study: Edge Growing v.s. Onion-Peeling. In this case study, we compare the application scenarios of $\mathsf{ASX}$-$\mathsf{OP}$ and $\mathsf{ASX}$-$\mathsf{I}$, as shown in Figure 9. We observe that $\mathsf{ASX}$-$\mathsf{OP}$, with onion peeling, cater users’ needs who “think slow” and seeks in-depth explanations that are counterfactual. Such larger and counterfactual explanations are more likely to affect $\mathsf{GNN}$ behaviors if removed [liu2021multi]. $\mathsf{ASX}$-$\mathsf{I}$, on the other hand, due to its containment of edges closer to targeted ones via edge insertions/growths, tends to produce – at its early stage – smaller, factual explanations as convincing evidences. This caters users who “think fast” and is good at quick generation of factual explanations [liu2021multi]. For example, in answering “why” node $v^{\prime}_{t}$ is recognized as a “Politician” site, $G_{\zeta}^{8}$ and $G_{\zeta}^{9}$ are among skyline explanations from $\mathsf{ASX}$-$\mathsf{I}$ at early stage, revealing the fact that $v^{\prime}_{t}$ frequently interact with TV shows, Governments, and Companies, covering its ego-networks as quick facts. $G_{\zeta}^{10}$ and $G_{\zeta}^{11}$, which are reported by $\mathsf{ASX}$-$\mathsf{OP}$ in its early stage via “onion-peeling”, are relatively larger counterfactual explanations that suggest interaction patterns with more complex structures. For example, $G_{\zeta}^{11}$ includes two-hop neighbors and a triangle, which indicates the importance of the corresponding company and TV show. Most importantly, another politician’s two-hop neighbor also shares these nodes, revealing the close relationship between the two politicians.